# backgroundcolor=, basicstyle=

### v1

@@ 1 Introduction
However, dual label datasets are rare and often small. They present interesting findings, but a more robust dataset is needed for full evaluation. We therefore also introduce a more comprehensive evaluation method for large-scale traditional datasets, such as the folktables datasets [13].
@@ 3.1 The Dataset Class
The currently implemented datasets, summarized in Tab. 2, consist of two types.
@@ 3.2 The FairnessMethod class
Ten fairness methods are currently implemented in ABCFair. An overview of these methods and their properties with respect to the comparability challenges in Sec. 2 is given in Tab .3.
@@
The Error Parity method [11] is the only postprocessing method currently implemented in ABCFair. It enjoys strong theoretical guarantees by creating a separate decision thresholds for each sensitive group such that they equalize a given fairness notion (up to a controllable bound).
@@ 3.3 The Evaluator class
The Evaluator class monitors the predictions during training, validation, and testing.
@@
Fairness is measured using the parity-based fairness definition in the fairret library. We refer to [6] for a full discussion, but summarize the methodology. First, the statistic $\gamma(q;h)$ for a relevant fairness notion is computed for every sensitive feature $q\in\{1,...,\left\lvert\mathcal{S}\right\rvert\}$, over the model scores $h(X)$.
We discuss results on two statistics $\gamma$ here, and report on five more in the Appendix.
@@
$\gamma(k;h)=\frac{\operatorname{\mathbb{E}}\left[S_{q}h(X)\right]}{\operatorname{\mathbb{E}}\left[S_{q}\right]},$ (1)
@@
$\gamma(k;h)=\frac{\operatorname{\mathbb{E}}\left[S_{k}h(X)Y\right]}{\operatorname{\mathbb{E}}\left[S_{k}Y\right]}$ (2)
@@
i.e. each $\gamma(q;h)$ is compared to the overall mean statistic $\overline{\gamma}(h)$, found by setting $S_{q}\equiv 1$. We take the maximal absolute value of their ratio.
@@ 4.2 Full result comparison on a real-world dataset
Key Finding 5: Whether the output consists of hard or soft scores has a significant impact on trade-offs. Again, we already argued this comparability problem in Sec. 2.4. Here, Exponentiated Gradient and Error Parity are both methods that do not optimize with soft model scores in mind, causing them to incur significant drops in performance in Tab. 4 and 5 for even the highest $k$ bounds (which correspond with the unfairness of the naive model). The Fairret $\textnormal{KL}_{\textnormal{proj}}$ method incurs no performance loss for all levels of fairness violation, as it does optimize for soft scores.
Conversely, Error Parity performs very well when its performance is measured as intended in Tab. 6: on hard, binary predictions.

### v2

@@ 1 Introduction
However, dual label datasets are rare and often small. They present interesting findings, but a more robust dataset is needed for full evaluation. We therefore also introduce a more comprehensive evaluation method for large-scale traditional datasets, such as the folktables datasets [15]. Which is used in many benchmarks, even for topics other than bias mitigation method [11, 20, 19, 21, 24, 1].
@@ 3.1 The Dataset Class
Tab. 2 provides details on the datasets available in the current implementation of the repository.
Experiment details.
@@ 3.2 The FairnessMethod class
Experiment details.
Ten fairness methods are currently implemented in ABCFair. Tab .3 provides an overview of these methods and their properties with respect to the comparability challenges.
@@ 3.2.3 Postprocessing
Experiment details.
Error Parity [13] is the only postprocessing method currently implemented in ABCFair. It enjoys strong theoretical guarantees by creating a separate decision thresholds for each sensitive group such that they equalize a given fairness notion (up to a controllable bound).
@@ 3.3 The Evaluator class
Implementation.
The Evaluator class monitors the predictions during training, validation, and testing. It computes the performance of the model for every prediction type, combination of sensitive attributes, fairness statistic, and performance measure.
Experiment details.
@@
Fairness is measured using the parity-based fairness definition in the fairret library. We refer to [7] for a full discussion, but summarize the methodology. First, the statistic $\gamma(q;h)$ for a relevant fairness notion is computed for every sensitive feature $q\in\{1,...,\left\lvert\mathcal{S}\right\rvert\}$, over the model scores $h(X)$. We discuss results on two statistics $\gamma$ here, and report on five more in the Appendix.
@@
$\gamma(q;h)=\frac{\operatorname{\mathbb{E}}\left[S_{q}h(X)\right]}{\operatorname{\mathbb{E}}\left[S_{q}\right]},$ (1)
@@
$\gamma(q;h)=\frac{\operatorname{\mathbb{E}}\left[S_{q}h(X)Y\right]}{\operatorname{\mathbb{E}}\left[S_{q}Y\right]}$ (2)
@@
i.e. each $\gamma(q;h)$ is compared to the overall mean statistic $\overline{\gamma}(h)$, found by setting $S_{q}\equiv 1$.
@@ 4.2 Full result comparison on a real-world dataset
Key Finding 5: Whether the output consists of hard or soft scores has a significant impact on trade-offs. Again, we already argued this comparability problem in Sec. 2.4. Here, Exponentiated Gradient and Error Parity are both methods that do not optimize with soft model scores in mind, causing them to incur significant drops in performance in Tab. 4 and 5 for even the highest $k$ bounds (which correspond with the unfairness of the naive model). The Fairret $\textnormal{KL}_{\textnormal{proj}}$ method incurs no performance loss for all levels of fairness violation, as it does optimize for soft scores. Conversely, Error Parity performs very well when its performance is measured as intended in Tab. 6: on hard, binary predictions.
@@ 5 Conclusions
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] The contribution and abstract both discuss how we introduce a new fairness benchmarking approach. We discuss the challenges, how we see to tackle these challenges, proposes a pipeline to accompany the approach and provide an example of benchmarking according to our approach.
    2. (b)
      Did you describe the limitations of your work? [Yes] The conclusions of the paper finish with discussing the limitations of the work.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] This work is meant to discuss and provide a solution to the difficulties faced in benchmarking fairness methods. Although a risk for fairness gerrymandering could exist, this risk is not a consequence of this work.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] The goal of this research is to reduce harmful impacts on society by the use of machine learning. We are fully transparent on the code and resources used for this project and ensure reproducibility of our results. All datasets used can be accessed at the location where the original authors have originally published them.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] The paper does not contain theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] The pipeline is made accessible by providing a link to its repository. This allows to fully reproduce our results.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] All information on data preprocessing and used hyperparameters are noted in the supplementary material.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] Fairness is a more complex notion which requires as to express the error as a confidence ellipse rather than a 1-dimensional error bar. More information about these confidence ellipses can be found in the appendix.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] The supplementary material contains information on the infrastructure that was used and the total compute time that was used for the experimental section.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] The original packages that implement these methods are cited and the papers that introduce these methods, if this information was provided by the package.
    2. (b)
      Did you mention the license of the assets? [Yes] Each code segment which makes use of code from an external source mentions its original packages and refers back to its license. All assets also contain proper reference to their source.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We provide a benchmark pipeline as a materialisation of our approach. This pipeline can be used for future benchmarking, a link to the repository containing all relevant information can be found in the supplementary material.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] All data used for this work came from publicly accessible sources. Information about consent can be found in the original sources of the datasets.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] The data used in this paper cannot be used to identify people as is discussed in the original sources of these datasets.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] We did not use crowdsourcing or conducted research with human experiments.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]